OpenStudy (anonymous):

Ive been stuck for hours.. Please help..? f(x)=2-2x Evaluate f(4) a 14 b -10 c -5 d 12 Which ordered pair is the turning point of the graph of f(x)=|x-4| a 4,0 b 0,4 c 0,-4 d -4,0

OpenStudy (b87lar):

I can help. Fornstarters though I'd like you tocheck the function - are u sure it is 2-2x, and not -2-2x?

OpenStudy (anonymous):

Oops it shouldnt be 2-2x it should be 2-3x.. i thought i fixed that

OpenStudy (anonymous):

input 4 into x

OpenStudy (anonymous):

i think you can continue from here

OpenStudy (b87lar):

Ok that makes more sense . Now one ofnthe four answers is possible.

OpenStudy (b87lar):

Do you knowhow to go about the second part?

OpenStudy (anonymous):

No, i do online schooling, and its a brand new teacher at teaching, and she barely went over the topic. & in our course there isnt much reading. :/ All it we really do is just quizzes after quizzes.. little time with the teacher

OpenStudy (anonymous):

|x-4| actually means -x+4=0

OpenStudy (b87lar):

Hmm I think it is not D. But I'd prefer if we came up with the solution together.

OpenStudy (anonymous):

i think you can take it from here

OpenStudy (anonymous):

i already got the solution on there

OpenStudy (anonymous):

x=4

OpenStudy (anonymous):

(x,y) Now what do u think the answer is if Y=0

OpenStudy (b87lar):

Would you be able to draw the graph of the function |x-4|

OpenStudy (anonymous):

i did that on my phone I think i actually just got the answer.. Would it be -10

OpenStudy (b87lar):

If you are still with the first part then yes it is -10.

OpenStudy (anonymous):

Am correct it is A

OpenStudy (anonymous):

for the second solution

OpenStudy (anonymous):

1.B 2.A

OpenStudy (mathmale):

I need to be absolutely sure of this: Does your "f(x)=2-2x" represent a linear function of the form y=mx+b, or did you mean "exponentiation," as in \[f(x)=2-2^x?\]